OpenStudy (gracygirl):

Medal and Fan Write the standard form of the line that passes through the point (-2, 4) and is parallel to x - 2y = 6. Type your answer in the box provided or use the upload option to submit your solution.

OpenStudy (gracygirl):

@Will.H

OpenStudy (3mar):

May I help?

OpenStudy (gracygirl):

Of course

OpenStudy (3mar):

Thank you.

OpenStudy (3mar):

Take that. Hope that helps

OpenStudy (will.h):

Parallel to X -2y= 6 To figure the slope of this set it in slope form -2y = -x + 6 Y = 1/2 X -3 That means slope should be 1/2 because the other equation will be parallel to this. (the share the slopes when they are parallel) Now use Y - y1 = m(x - x1) Where m is the slope And we have point (-2,4) to use So substitute Y - 4 = 1/2(x + 2) Y - 4 =( 1/2)x + 1 Standard form means x and y on the same side so make that happen Y - (1/2)x = 5 That's the equation hope that helps

OpenStudy (gracygirl):

Thank you both!

OpenStudy (3mar):

For m (the slope of the nine): all parallel lines have the same slope. So it'd better to find out the slope of the parallel line then equate it to the slope of the required line.

OpenStudy (will.h):

My pleasure.

OpenStudy (3mar):

You are welcome! Any Help... Any time.